A lottery claims its grand prize is %15 million, payable over 5 years at $3,000,000 per year. If the first payment is made immediately, what is the grand prize really worth? Use an interest rate of 4%.

The real value of the grand prize is ?
Answer #1

ANSWER:

Annual amount = $3,000,000

n = 4 years (first payment made in year 0)

i = 4%

Present worth = Annual amount in year 0 + annual amount from year 1 to year 4(p/a,i,n)

Present worth = 3,000,000 + 3,000,000(p/a,4%,4)

Present worth = 3,000,000 + 3,000,000 * 3.63

Present worth = 3,000,000 + 10,890,000

Present worth = 13,890,000

so the real value is $13,890,000